Hi, I am really stuck on this question. Can someone help me please??
A secretary periodically checks to see how many of the three lines into the office are busy. her finding for one week were the following.

No. lines busy Frequency
0 20
1 65
2 25
3 10
Let the random variable X be the number of lines busy. Find the probability distribution of X.

Answer by edjones(8007)

20+65+25+10=120
.
0 20/120=.1667
1 65/120=.5417
2 25/120=.2083
3 10/120=.0833
-----------------
========>1.0000
.
Ed